# A Low Cost Relativistic Algebraic Diagrammatic Construction Method Based on Cholesky Decomposition and Frozen Natural Spinors for Electronic Ionization, Attachment and Excitation Energy Problem

### v1

@@ 1 Introduction
The equation-of-motion coupled cluster (EOM-CC)[krylov2008equation, nooijen1992coupled, stanton1993equation, rowe1968equations] method has become one of the most widely used direct difference-based wavefunction methods for computing excitation energy (EE), electron attachment (EA), and ionization potential (IP) energy. The linear-response coupled cluster (LR-CC)[koch1990excitation, sekino1984linear] method gives identical excitation energies, although it is formulated from a different theoretical perspective than the EOM-CC approaches. However, in both cases, the non-Hermitian nature of the effective Hamiltonian within a basis of the biorthogonal CC framework makes the eigenvalue problems prone to complex solutions. [thomas2021complex] Moreover, the computation of transition properties in the standard CC method is challenging, as it requires the determination of both left and right eigenvectors. The unitary coupled cluster (UCC) approach offers a Hermitian formulation of CC methods and helps to address these challenges.[liu2018unitary, liu2021unitary] However, the non-termination Baker-Campbell-Hausdorff (BCH) expansion of the similarity-transformed Hamiltonian poses significant challenges for practical implementation of the UCC method. To address this issue, commutator-based truncation schemes within UCC theory has attracted considerable attention for due to their implication in quantum computing. However, they are shown to be advantageous even on classical computers.[majee2024reduced, majee2025relativistic]
@@
In contrast of these iterative nature of CC- based method the Algebraic Diagrammatic Construction (ADC)[schirmer2004intermediate, mertins1996algebraic, schirmer1991closed, schirmer1982beyond, von1984computational, dreuw2015algebraic, dempwolff2019efficient], scheme has attracted considerable attention due to its low-cost, non-iterative treatment of the ground state wave-function based on Møller–Plesset perturbation theory (MPn), as well as its Hermitian formulation, which provides distinct advantages in the calculation of excited state properties. The ADC method was first developed using the propagator formalism[schirmer1982beyond, trofimov1999consistent]. The propagator-based ADC methods were successfully applied to compute electron-ionization and attachment spectra[fetter2012quantum, cederbaum1975one, cederbaum1977theoretical, schirmer1978two]. In modern implementations of ADC, the intermediate-state representation (ISR) has emerged as the preferred approach, largely due to its transparent formalism and its provision of direct access to the excited-state wavefunction.[schirmer1991closed] An extension of the fourth-order ADC (ADC(4)) scheme within the ISR formalism for EE, IP, and EA has been proposed by Dreuw and co-workers.[leitner2022fourth, leitner2024fourth] In recent years, ADC methods have been extensively employed for energy and property calculations, particularly in lighter elements[leitner2024fourth, schirmer1983new, angonoa1987theoretical, dreuw2015algebraic, trofimov1995efficient, trofimov1997polarization, trofimov1999consistent, schirmer2004intermediate, trofimov2005molecular]. For atoms and molecules containing heavy elements, the inclusion of relativistic effects becomes essential. One of the most straightforward ways of introducing relativistic effects in quantum chemical calculations is to use a four-component Dirac Coulomb Hamiltonian.[dyall2007introduction] Pernpointer et al. have implemented the second-order and the extended second-order versions of ADC methods based on the four-component relativistic Hamiltonian, demonstrating its capability to compute excitation energies and transition dipole moments. [pernpointner2014relativistic, pernpointner2018four] In addition, non-Dyson-based second and third-order ADC schemes have been developed within the relativistic framework to compute valence ionization energies and investigate the electron decay process[pernpointner2004one, pernpointner2004one, pernpointner2010four, pernpointner2005effect]. Although the second-order ADC (ADC(2)) performs can give qualitative agreement, it typically lacks the quantitative accuracy. To achieve more reliable results, it is necessary to go beyond second-order and include at least third-order contributions in the perturbative expansion. To address this, we have recently developed a four-component(4c) ADC(3) method for calculating IP, EA, EE, and excited state properties of atoms and molecules containing heavy elements[chakraborty2025relativistic]. However, performing 4c-ADC(3) calculations remains a significant challenge due to their extremely high computational cost, often up to 32 times greater than that of a non-relativistic method[dyall2007introduction]. As a result, such calculations are typically limited to atoms and small molecules using a moderate-sized basis set. To mitigate this limitation, numerous approaches have been developed in the literature to lower the computational cost in the relativistic electronic structure calculation. One of the most effective strategies is the use of two-component relativistic theories[dyall2007introduction, hess1986relativistic, van1996relativistic, dyall1997interfacing, nakajima1999new, liu2009exact, saue2011relativistic]. Among the various flavours of two-component methods available, the exact two-component (X2C) theory based on atomic mean field (AMF) spin-orbit integrals (X2CAMF) has emerged as one of the most attractive ones[liu2018atomic, zhang2022atomic, knecht2022exact]. It can be further combined with the cholesky decomposition(CD) treatment of the integrals[helmich2019relativistic, banerjee2023relativistic, uhlirova2024cholesky]. Sokolov and co-workers[Sokolov] have recently reported a single-reference and multi-reference ADC method for ionization and electron attachment problems based on a one-electron variant of the X2C method (X2C-1e). In addition to the simplification offered by the two-component framework, leveraging massively parallel programs can accelerate the floating-point operations of relativisitic electron correlation method.[pototschnig2021implementation] Alternatively, one can use a more compact set of spinor space in the form of frozen natural spinors (FNS), which allows one to perform effective truncation of the virtual space with systematic controllable accuracy. The FNS are generally constructed from MP2 calculations, and the efficiency of the MP2 based FNS has been demonstrated in relativistic four-component ground-state coupled cluster (CC)[chamoli2022reduced] and unitary coupled cluster (UCC)[majee2024reduced] calculations. This approach has also been extended to core and valence ionization problems[surjuse2022low] using the equation of motion coupled cluster approach(EOM-CC). Gomes and co-workers[yuan2022assessing] have reported a similar natural spinor-based strategy for the ground-state CC method. To further reduce the computational cost, the FNS method can be combined with the CD decomposition-based X2CAMF-CC method[chamoli2025frozen]. Gomes and coworkers[yuan2022assessing] have demonstrated that, in contrast to ground state energy calculation, standard MP2-based natural spinors fail to describe the excited states due to their inability to capture the corresponding electronic distribution. It has been recently demonstrated that a state-specific approach can mitigate this problem[mukhopadhyay2025reduced]. In this work, we present the theory, implementation, and benchmarking of a low-cost relativistic ADC(3) method for ionized, electron-attached, and excited states using the CD-X2CAMF Hamiltonian and frozen natural spinor approximation.
@@ 2.1 Relativistic Algebraic Diagrammatic Construction Theory
The algebraic–diagrammatic construction (ADC) method offers a a systematic scheme for the calculations of generalized excitation energies[schirmer1982beyond, dreuw2015algebraic, schirmer2018many, schirmer1991closed, oddershede1978polarization, ODDERSHEDE198433, oddershede1987propagator, schirmer1983new, ortiz2013electron] within the framework of propagator theory. The type of spectroscopic information obtained depends on the propagator under consideration. For instance, employing ADC in connection with the polarization propagator provides access to electronic excitation spectra,[schirmer1982beyond, dreuw2015algebraic, schirmer2018many, trofimov1995efficient, trofimov1997polarization, trofimov1997polarizationII], whereas the formulation based on the electron propagator enables the calculation of ionization potentials and electron affinities.[schirmer2018many, schirmer1983new, von1984computational, schirmer1989green, schirmer1998non, trofimov2005molecular]
@@ 2.4 Cholesky decomposition
In this implementation, four-external () and three-external () integrals are neither precomputed nor stored, but generated on the fly. All other integrals are explicitly constructed and stored in the truncated FNS basis. More details regarding the FNS and CD-based X2CAMF implementation of relativistic coupled-cluster methods are available in Ref. [chamoli2025frozen].
@@ 2.5 Frozen natural spinors
where represents the MP2-level virtual–virtual density block and is the electron-attached/excited-state contribution from EA/EE-ADC(2). The zeroth-order intermediate state representation[schirmer2004intermediate] with ADC(2) eigen vectors were used to obtain excited-state densities. The programmable expression for the state-specific density are provided in the supporting information. Inserting Eq. (47) into Eq. (42) and following Eqs. (42–46) produces the state-specific frozen natural spinor basis (SS-FNS) for each excited state, in which the ADC(3) calculations are subsequently performed. A schematic representation of the algorithm used for EA-ADC(3) and EE-ADC(3) has been presented in Figure1the supporting information.
@@ 3.1.1 Choice of truncation threshold
The accuracy of the FNS based method depends upon the occupation threshold used for truncation. To get an appropriate FNS threshold, we have studied the convergence of IP values with respect to the FNS threshold for the first IP value of IBr(See Figure 2). The uncontracted dyall.v2z basis set has been used for the calculations. The CD threshold has been kept at following reference [chamoli2025reduced]. It can be seen that the result almost converges at the threshold of . The error is even smaller at the threshold of . Therefore, we have used the CD threshold and FNS threshold of for the subsequent calculation, which is similar to the LOOSEFNS threshold in reference[chamoli2025reduced].
@@ 3.2.4 Transition properties of Xe atom
Table 5 reports the excitation energies (EE) and transition dipole moments (TDMs) for selected excited states of the xenon atom, calculated using SS-FNS-EE-ADC(3) and 4c-EE-ADC(3), and compared against experimental data. The calculations employed the d-aug-Dyall.ae3z basis set, with the frozen-core approximation and a virtual spinor energy cutoff of 361 Eh. This setup yields a correlation space of 18 occupied and 364 virtual spinors. The same basis set and the correlation space is used as in our previous 4c-EE-ADC(3) implementation.[chakraborty2025relativistic]. The SS-FNS-EE-ADC(3) closely reproduces both the EEs and TDMs obtained from 4c-EE-ADC(3). The deviations between SS-FNS and 4c are within 0.03 eV for EEs and less than 0.005 a.u. for TDMs, indicating that in the SS-FNS framework, the truncated virtual space is sufficient to capture the essential correlation and relativistic effects for these transitions. Interestingly, with a truncation threshold of , approximately 56% of the 364 virtual spinors are excluded, leaving only 158 active virtual spinors for the SS-FNS-EE-ADC(3) calculations. However, the performance of SS-FNS-EE-ADC(3) is mixed compared to the experimental results. The calculated transition moment shows good agreement with the experimental results except for the states. However, the excitation energy values are significantly underestimated as compared to the experiments for all states. The performance is similar to that of the canonical 4c-ADC(3) method. The excitation energy values can be improved by using the scaled matrix SS-FNS-EE-ADC[(2)+x(3)] method. However, the agreement with experiment for transition moment significantly deteriorates on using SS-FNS-EE-ADC[(2)+x(3)]. The trend is consistent with that observed by Dreuw and co-workers [bauerExploringAccuracyUsefulness2022] for the non-relativistic case.
@@ 3.5 Application to medium and large systems
For showcasing the performance of our SS-FNS-EE-ADC(3) implementation, we have considered complex, and the excitation energy corresponding to the lowest excited state has been calculated. The geometry of the hexa-aqua triiodide complex has been obtained from Ref.[mukhopadhyay2025reduced] where they have used the same system for demonstrating the performance of SS-FNS-EE-EOM-CCSD method. The dyall.av3z basis set has been used for the I atoms, and the uncontracted version of the cc-pVDZ basis set has been used for the O and H atoms. The hexa-aqua triiodide complex consists of 21 atoms with 220 electrons. The aforementioned basis set uses 1394 virtual spinors on a canonical basis and a Cholesky threshold of has been used resulting in 2580 Cholesky vectors. The SS-FNS-EE-ADC(3) calculation has been performed in a dedicated workstation with two Intel(R) Xeon(R) Gold 5315Y CPUs @ 3.20 GHz and 2 TB of total RAM. After applying frozen core approximation and using an FNS threshold of , the correlation space in the truncated basis contains 70 occupied and 517 virtual spinors. The total time taken for this calculation is 3 days, 18 hours, 25 minutes, and 14 seconds, whereas with a similar computational setting, the SS-FNS-EE-EOM-CCSD took 5 days, 1 hour, 41 minutes. The total time taken in the ADC(3) part in the truncated basis is 2 days, 8 hours, 40 minutes, compared to 3 days, 16 hours, 32 minutes taken by the EOM-CCSD version. The excitation energy obtained for this complex of the state is 2.16 eV.

### v2

@@ 1 Introduction
The equation-of-motion coupled cluster (EOM-CC)[krylov2008equation, nooijen1992coupled, stanton1993equation, rowe1968equations] method has become one of the most widely used direct difference-based wavefunction methods for computing excitation energy (EE), electron attachment (EA), and ionization potential (IP) energy. The linear-response coupled cluster (LR-CC)[koch1990excitation, sekino1984linear] method gives identical excitation energies, although it is formulated from a different theoretical perspective than the EOM-CC approaches. However, in both cases, the non-Hermitian nature of the effective Hamiltonian within a basis of the biorthogonal CC framework makes the eigenvalue problems prone to complex solutions. [thomas2021complex] Moreover, the computation of transition properties in the standard CC method is challenging, as it requires the determination of both left and right eigenvectors. The unitary coupled cluster (UCC) approach offers a Hermitian formulation of CC methods and helps to address these challenges.[liu2018unitary, liu2021unitary] However, the non-termination Baker-Campbell-Hausdorff (BCH) expansion of the similarity-transformed Hamiltonian poses significant challenges for practical implementation of the UCC method. To address this issue, commutator-based truncation schemes within UCC theory have garnered considerable attention due to their implications in quantum computing. However, they are shown to be advantageous even on classical computers.[majee2024reduced, majee2025relativistic]
@@
In contrast of these iterative nature of CC- based method the Algebraic Diagrammatic Construction (ADC)[schirmer2004intermediate, mertins1996algebraic, schirmer1991closed, schirmer1982beyond, von1984computational, dreuw2015algebraic, dempwolff2019efficient], scheme has attracted considerable attention due to its low-cost, non-iterative treatment of the ground state wave-function based on Møller–Plesset perturbation theory (MPn), as well as its Hermitian formulation, which provides distinct advantages in the calculation of excited state properties. The ADC method was first developed using the propagator formalism[schirmer1982beyond, trofimov1999consistent]. The propagator-based ADC methods were successfully applied to compute electron-ionization and attachment spectra[fetter2012quantum, cederbaum1975one, cederbaum1977theoretical, schirmer1978two]. In modern implementations of ADC, the intermediate-state representation (ISR) has emerged as the preferred approach, largely due to its transparent formalism and its provision of direct access to the excited-state wavefunction.[schirmer1991closed] An extension of the fourth-order ADC (ADC(4)) scheme within the ISR formalism for EE, IP, and EA has been proposed by Dreuw and co-workers.[leitner2022fourth, leitner2024fourth] In recent years, ADC methods have been extensively employed for energy and property calculations, particularly in lighter elements[leitner2024fourth, schirmer1983new, angonoa1987theoretical, dreuw2015algebraic, trofimov1995efficient, trofimov1997polarization, trofimov1999consistent, schirmer2004intermediate, trofimov2005molecular]. For atoms and molecules containing heavy elements, the inclusion of relativistic effects becomes essential. One of the most straightforward ways of introducing relativistic effects in quantum chemical calculations is to use a four-component Dirac Coulomb Hamiltonian.[dyall2007introduction] Pernpointer et al. have implemented the second-order and the extended second-order versions of ADC methods based on the four-component relativistic Hamiltonian, demonstrating its capability to compute excitation energies and transition dipole moments. [pernpointner2014relativistic, pernpointner2018four] In addition, non-Dyson-based second and third-order ADC schemes have been developed within the relativistic framework to compute valence ionization energies and investigate the electron decay process[pernpointner2004one, pernpointner2004one, pernpointner2010four, pernpointner2005effect]. Although the second-order ADC (ADC(2)) can give qualitative agreement, it typically lacks quantitative accuracy. To achieve more reliable results, it is necessary to go beyond second-order and include at least third-order contributions in the perturbative expansion. To address this, we have recently developed a four-component(4c) ADC(3) method for calculating IP, EA, EE, and excited state properties of atoms and molecules containing heavy elements[chakraborty2025relativistic]. However, performing 4c-ADC(3) calculations remains a significant challenge due to their extremely high computational cost, often up to 32 times greater than that of a non-relativistic method[dyall2007introduction]. As a result, such calculations are typically limited to atoms and small molecules using a moderate-sized basis set. To mitigate this limitation, numerous approaches have been developed in the literature to lower the computational cost in the relativistic electronic structure calculation. One of the most effective strategies is the use of two-component relativistic theories[dyall2007introduction, hess1986relativistic, van1996relativistic, dyall1997interfacing, nakajima1999new, liu2009exact, saue2011relativistic]. Among the various flavours of two-component methods available, the exact two-component (X2C) theory based on atomic mean field (AMF) spin-orbit integrals (X2CAMF) has emerged as one of the most attractive ones[liu2018atomic, zhang2022atomic, knecht2022exact]. It can be further combined with the cholesky decomposition(CD) treatment of the integrals[helmich2019relativistic, banerjee2023relativistic, uhlirova2024cholesky]. Sokolov and co-workers[Sokolov] have recently reported a single-reference and multi-reference ADC method for ionization and electron attachment problems based on a one-electron variant of the X2C method (X2C-1e). In addition to the simplification offered by the two-component framework, leveraging massively parallel programs can accelerate the floating-point operations of relativisitic electron correlation method.[pototschnig2021implementation] Alternatively, one can use a more compact set of spinor space in the form of frozen natural spinors (FNS), which allows one to perform effective truncation of the virtual space with systematic controllable accuracy. The FNS are generally constructed from MP2 calculations, and the efficiency of the MP2-based FNS has been demonstrated in relativistic four-component ground-state coupled cluster (CC)[chamoli2022reduced] and unitary coupled cluster (UCC)[majee2024reduced] calculations. This approach has also been extended to core and valence ionization problems[surjuse2022low] using the equation of motion coupled cluster approach(EOM-CC). Gomes and co-workers[yuan2022assessing] have reported a similar natural spinor-based strategy for the ground-state CC method. To further reduce the computational cost, the FNS method can be combined with the CD decomposition-based X2CAMF-CC method[chamoli2025frozen]. Gomes and coworkers[yuan2022assessing] have demonstrated that, in contrast to ground state energy calculations, standard MP2-based natural spinors fail to describe excited states due to their inability to capture the corresponding electronic distribution. It has been recently demonstrated that a state-specific approach can mitigate this problem[mukhopadhyay2025reduced]. In this work, we present the theory, implementation, and benchmarking of a low-cost relativistic ADC(3) method for ionized, electron-attached, and excited states using the CD-X2CAMF Hamiltonian and frozen natural spinor approximation.
@@ 2.1 Relativistic Algebraic Diagrammatic Construction Theory
The algebraic–diagrammatic construction (ADC) method offers a systematic scheme for the calculations of generalized excitation energies[schirmer1982beyond, dreuw2015algebraic, schirmer2018many, schirmer1991closed, oddershede1978polarization, ODDERSHEDE198433, oddershede1987propagator, schirmer1983new, ortiz2013electron] within the framework of propagator theory. The type of spectroscopic information obtained depends on the propagator under consideration. For instance, employing ADC in connection with the polarization propagator provides access to electronic excitation spectra,[schirmer1982beyond, dreuw2015algebraic, schirmer2018many, trofimov1995efficient, trofimov1997polarization, trofimov1997polarizationII], whereas the formulation based on the electron propagator enables the calculation of ionization potentials and electron affinities.[schirmer2018many, schirmer1983new, von1984computational, schirmer1989green, schirmer1998non, trofimov2005molecular]
@@ 2.4 Cholesky decomposition
Although we are using the one-step procedure in the current work, but it is worth noting that an efficient two-step algorithm for the Cholesky decomposition of the electron-repulsion integrals has also been developed, in which only the elements of the Cholesky basis are generated during the pivoting procedure.[aquilante2011cholesky, folkestad2019efficient, zhang2021toward]. In this implementation, four-external () and three-external () integrals are neither precomputed nor stored, but generated on the fly. All other integrals are explicitly constructed and stored in the truncated FNS basis. More details regarding the FNS and CD-based X2CAMF implementation of relativistic coupled-cluster methods are available in Ref. [chamoli2025frozen].
@@ 2.5 Frozen natural spinors
where represents the MP2-level virtual–virtual density block and is the electron-attached/excited-state contribution from EA/EE-ADC(2). The zeroth-order intermediate state representation[schirmer2004intermediate] with ADC(2) eigen vectors were used to obtain excited-state densities. The programmable expression for the state-specific density are provided in the supporting information. Inserting Eq. (47) into Eq. (42) and following Eqs. (42–46) produces the state-specific frozen natural spinor basis (SS-FNS) for each excited state, in which the ADC(3) calculations are subsequently performed. A schematic representation of the algorithm used for EA-ADC(3) and EE-ADC(3) has been presented in Figure1 in the supporting information.
@@ 3.1.1 Choice of truncation threshold
The accuracy of the FNS-based method depends upon the occupation threshold used for truncation. To get an appropriate FNS threshold, we have studied the convergence of IP values with respect to the FNS threshold for the first IP value of IBr(See Figure 2). The uncontracted dyall.v2z basis set has been used for the calculations. The CD threshold has been kept at following reference [chamoli2025reduced]. It can be seen that the result almost converges at the threshold of . The error is even smaller at the threshold of . Therefore, we have used the CD threshold and FNS threshold of for the subsequent calculation, which is similar to the LOOSEFNS threshold in reference[chamoli2025reduced].
@@ 3.2.4 Transition properties of Xe atom
Table 5 reports the excitation energies (EE) and transition dipole moments (TDMs) for selected excited states of the xenon atom, calculated using SS-FNS-EE-ADC(3) and 4c-EE-ADC(3), and compared against experimental data. The calculations employed the d-aug-Dyall.ae3z basis set, with the frozen-core approximation and a virtual spinor energy cutoff of 361 Eh. This setup yields a correlation space of 18 occupied and 364 virtual spinors. The same basis set and the correlation space are used as in our previous 4c-EE-ADC(3) implementation.[chakraborty2025relativistic]. The SS-FNS-EE-ADC(3) closely reproduces both the EEs and TDMs obtained from 4c-EE-ADC(3). The deviations between SS-FNS and 4c are within 0.03 eV for EEs and less than 0.005 a.u. for TDMs, indicating that in the SS-FNS framework, the truncated virtual space is sufficient to capture the essential correlation and relativistic effects for these transitions. Interestingly, with a truncation threshold of , approximately 56% of the 364 virtual spinors are excluded, leaving only 158 active virtual spinors for the SS-FNS-EE-ADC(3) calculations. However, the performance of SS-FNS-EE-ADC(3) is mixed compared to the experimental results. The calculated transition moment shows good agreement with the experimental results except for the states. However, the excitation energy values are significantly underestimated as compared to the experiments for all states. The performance is similar to that of the canonical 4c-ADC(3) method. The excitation energy values can be improved by using the scaled matrix SS-FNS-EE-ADC[(2)+x(3)] method. However, the agreement with experiment for transition moment significantly deteriorates on using SS-FNS-EE-ADC[(2)+x(3)]. The trend is consistent with that observed by Dreuw and co-workers [bauerExploringAccuracyUsefulness2022] for the non-relativistic case.
@@ 3.5 Application to medium and large systems
For showcasing the performance of our SS-FNS-EE-ADC(3) implementation, we have considered complex, and the excitation energy corresponding to the lowest excited state has been calculated. The geometry of the hexa-aqua triiodide complex has been obtained from Ref.[mukhopadhyay2025reduced] where they have used the same system for demonstrating the performance of the SS-FNS-EE-EOM-CCSD method. The dyall.av3z basis set has been used for the I atoms, and the uncontracted version of the cc-pVDZ basis set has been used for the O and H atoms. The hexa-aqua triiodide complex consists of 21 atoms with 220 electrons. The aforementioned basis set utilizes 1394 virtual spinors on a canonical basis, and a Cholesky threshold of has been employed, resulting in 2580 Cholesky vectors. The SS-FNS-EE-ADC(3) calculation has been performed in a dedicated workstation with two Intel(R) Xeon(R) Gold 5315Y CPUs @ 3.20 GHz and 2 TB of total RAM. After applying frozen core approximation and using an FNS threshold of , the correlation space in the truncated basis contains 70 occupied and 517 virtual spinors. The total time taken for this calculation is 3 days, 18 hours, 25 minutes, and 14 seconds, whereas with a similar computational setting, the SS-FNS-EE-EOM-CCSD took 5 days, 1 hour, 41 minutes. The total time taken in the ADC(3) part in the truncated basis is 2 days, 8 hours, 40 minutes, compared to 3 days, 16 hours, 32 minutes taken by the EOM-CCSD version. The excitation energy obtained for this complex of the state is 2.16 eV.